OpenStudy (anonymous):

Show the work to solve |5x – 2| greater than or equal to 8 Describe the graph of the solution in words

OpenStudy (anonymous):

Keep the definition of the absolute value in mind: \[|x| = x\;\text{if}\; x \le 0\;\;\;|x| = -x\;\text{if}\; x \ge 0\]. So we distinguish two cases to be able to continue our discussion: 1) 5x - 2 > 0 => |5x-2| = 5x-2 Out of 5x - 2 > 0 follows 5x > 2 x > 2/5 So by supposing that 5x-2>0 it follows that x is greater than 2/5. Now we look at |5x - 2| = 5x - 2 >= 8. 5x - 2 >= 8 | +2 5x >= 10 | :2 x >= 2 So all x greater or equal to 2 satisfy our inequality. Now we look at the second case 2) 5x - 2 < 0 equivalent to x < 2/5. From that follows BY DEFINITION (see the above) 5x - 2 < 0 => |5x - 2| = -(5x -2) = -5x + 2 >= 8 -5x + 2 >= 8 | - 2 -5x >= 6 |:(-5) when dividing by a negative number, the inequality turns! x <= -6/5 So, our solutions are all x smaller than or equal to -6/5 and all x greater than or equal to 2. So, on the real line It's just anything left of or equal to -6/5 and anything right of or equal to 2

OpenStudy (anonymous):

thank you